A boat on a river travels 50miles downstream in only 5 hours. It takes the same boat 9 hours to travel 18miles upstream. What are the speed of the boat and the speed of the current?
Answer by mananth(16946)

boat speed= x mph
current speed= y mph
with current speed (x+y)
against current speed (x-y)
d= 50 downstream
50 / (x+y) = 5
divide by 5
10 / (x+y) = 1
(x+y) = 10 ............1
d=18
18 /(x-y)= 9 against wind
divide by 9
2 /(x-y) =
x-y=2 .............2
add up (1) & (2)
2 x = 12
/ 2
x= 6 mph speed of boat in still water
plug value of x in (1)
we get y= 4 mph speed of current